OpenStudy (anonymous):

Integrate: (looking for substitution or substitution by parts method, I think I \(\pm\) sign went wrong somewhere in my work... \[\huge \int\limits_{0}^{\pi} \frac{\cos(8θ)}{e^{θ}} d\theta\]

OpenStudy (anonymous):

It keeps getting uglier with each subsequent substitution so I'm guess I have something not canceling out or there's some trick I'm not seeing at this time. ;-)

OpenStudy (turingtest):

do you know the\[\int e^x\sin x dx\]integral trick? it's pretty well-known, and is the basis of how I would approach this problem

OpenStudy (anonymous):

This is what I was getting for what you just wrote u = e^x , du = e^x dx dv = sin(x) dx , v = -cos(x) \[\int e^x\sin x dx =-e^xcosx-\int e^x(-\cos x) dx\]

OpenStudy (anonymous):

If it was just \(\int\)xsinx dx it would be... sin(x)-xcos(x)+C yes?

OpenStudy (turingtest):

yes to you last post and continue with the other if you can\[\int e^x\sin x dx =-e^xcosx-\int e^x(-\cos x) dx=?\]

OpenStudy (anonymous):

Ah a second set of substitutions gives you \(e^x \sin x dx\) again... Hmm... Anti-derivative of +cos(x) \(\rightarrow\) +sin(x), can move the - out of the integral as -(-1) = +1 We get the same thing twice, which if it works like any other variable we can double it so... \[\frac{1}{2} e^xsin(x)-e^xcos(x)+C\] Still good?

OpenStudy (turingtest):

yes, but you missed the parentheses\[\frac12(e^2\sin x-e^x\cos x)+C\]so you are familiar with the "feedback" trick then?

OpenStudy (turingtest):

(where we wind up with the same integral we started with, and use algebra to solve for it)

OpenStudy (turingtest):

or maybe you missed the parentheses for lack of understanding something? let me type it out so you can see...

OpenStudy (turingtest):

now it should be pretty much the same story with your integral, with uglier numbers

OpenStudy (turingtest):

\[\int e^x\sin x dx =-e^xcosx-\int e^x(-\cos x) dx\]\[\int e^x\sin x dx =-e^x\cos x+e^x\sin x-\int e^x\sin xdx\]adding\(\large\int e^x\sin xdx\) to both sides gives\[2\int e^x\sin xdx=e^x(\sin x-\cos x)+C\]which we can solve for the integral (note that C/2 is still just some unknown constant, so we're gonna still call it C again after division)\[\int e^x\sin xdx={e^x(\sin x-\cos x)\over2}+C\]

OpenStudy (turingtest):

as I said earlier, same process, uglier numbers in your prob

OpenStudy (anonymous):

Following what you've written (which is quite clear, ty)... and trying it with what I have to do (and yes you're right I forgot the parenthesis) I have: \[[\frac{8}{65} e^{-\theta } \sin(8 \theta )-\frac{1}{65} e^{-\theta} cos(8 \theta)+C]_0^{\pi}\]

OpenStudy (anonymous):

Ty for your help again Turing :D

OpenStudy (turingtest):

welcome :)